OpenStudy (anonymous):

A loop of string is formed into a rectangle. If the width of the rectangle is 12 inches the height is 6 inches. If The width is 10 inches the height is 8 inches let w= width and h= height. If the width is 15 inches what is the height .??? PLEASE HELP ME

OpenStudy (anonymous):

solve the perimeter for either of the first 2 rectangles... 12(2)+6(2)=36 or 10(2)+8(2)=36 so, the formula for perimeter P is 2w+2h=P plug in given w 15(2)+2h=36...has to equal 36 solve for h

OpenStudy (anonymous):

How would you solve for H ?

OpenStudy (anonymous):

you want to have h just on one side and everything else on the other side of the equal sign. so first, look at the left side of the equal sign. it is 15(2)+2h or 30+2h the next step is to look for any terms that have no h which in this case, would be 30. you want to get rid of it so subtract 30 from this equation. what's done on one side of the equation must be done to the other side as well. so now you have 2h=36-30 or 2h=6 now looking at the left side of the equal sign again, we see that there's only one term left. to get rid of the 2, you have to divide by 2 since 2h is 2 times h and opposite of multiply is divide. again, what's done on one side must be done on the other side. so now you have h=6/2 or h=3.

OpenStudy (anonymous):

give two order pairs for the situation.

OpenStudy (anonymous):

didn't you just need the height? i don't understand what you're asking ^

OpenStudy (anonymous):

i dont honestly know this is my math question it says give to orderd pairs for the sutuastion